A doctor opens a 1.40 m wide door by pushing on it with a force of 52.5 N directed perpendicular to its surface.



(a)

What magnitude torque (in N · m) does she apply about an axis through the hinges if the force is applied at the center of the door?

__ N · m


(b)

What magnitude torque (in N · m) does she apply at the edge farthest from the hinges?

 __ N · m


1
Expert's answer
2021-10-11T08:55:34-0400

(a) By the definition of the torque, we have:


"\tau=Frsin\theta=52.5\ N\cdot\dfrac{1.40\ m}{2}\cdot sin90^{\circ}=36.75\ N\cdot m."

(b) By the definition of the torque, we have:


"\tau=Frsin\theta=52.5\ N\cdot1.40\ m\cdot sin90^{\circ}=73.5\ N\cdot m."